the length of a rectangle is three inches more than the width. The area of the rectangle is 418 inches. find the width of the rectangle
Answer by mananth(16946)

width = w
length = w+3
Area = L*W
w(w+3)=418
w^2+3w=418
w^2+3w-418=0
w^2+22w-19w-416=0
w(w+22)-19(w+22)=0
(w+22)(w-19)=0
w=19 inches which is positive
Length = 19+3=22 inches